OpenStudy (anonymous):

how to solve for this polynomial 4x+5y+9y-2y=13x+13y

OpenStudy (anonymous):

get all the x's on one side, and then all the y's on the other, then factor out the y's and x's. then divide through to solve for either y or x

OpenStudy (anonymous):

can you show how to do it?

OpenStudy (anonymous):

get anything multiplied by an x on one side, and anything multiplied by a y on the other, then type what you get

OpenStudy (anonymous):

so for instance, to move the 4x over to the right, i would subtract 4x from both sides

OpenStudy (anonymous):

i dnt understand

OpenStudy (anonymous):

so at the moment you have "some equation = some other equation" the equals sign means that both sides have the same value. so adding something to one side means you have to add to the other to keep the same value

OpenStudy (anonymous):

you need to move the 4x across. so subtract 4x from the right, and from the left

OpenStudy (anonymous):

can you show step by step I am really confused

OpenStudy (anonymous):

4x+ 5x + 9y - 2y = 13x + 13y is your equation

OpenStudy (anonymous):

k

OpenStudy (anonymous):

so if i subtract 4x from both sides i get 5x + 9y - 2y = 13x + 13y - 4x

OpenStudy (anonymous):

i have to do it to both sides at the same time, so it stays balanced.

OpenStudy (anonymous):

do you see why?

OpenStudy (anonymous):

does this make sense so far

OpenStudy (anonymous):

yes so what do i do next?

OpenStudy (anonymous):

so now you need to get rid off the 13y on the right hand side. Do that and type wwhat you get

OpenStudy (anonymous):

13y-5x+9y-2y=13x-4x

OpenStudy (anonymous):

yeah awesome! i think you misstyped the 5y for 5x but thats great

OpenStudy (anonymous):

so we have 13y-5y+9y-2y=13x-4x

OpenStudy (anonymous):

now we use the sneaky trick of factoring in order to clean up all those untidy y's

OpenStudy (anonymous):

y(13-5+9-2) = 13x - 4x

OpenStudy (anonymous):

can you see what ive done? ive "pulled" the y out from each bit in the equation.

OpenStudy (anonymous):

and so y mutiplies by the whole thing at once, instead of each bit in turn

OpenStudy (anonymous):

get it?

OpenStudy (anonymous):

so what do i do next?

OpenStudy (anonymous):

okay now you do the same with the x side, and type what you get

OpenStudy (anonymous):

y(13-5+9-2)=x(13-4)

OpenStudy (anonymous):

yeah awesome which is y15 = x9

OpenStudy (anonymous):

are you solving this for y or for x?

OpenStudy (anonymous):

both

OpenStudy (anonymous):

you have to chose to solve it for y in terms of x, or x in terms of y

OpenStudy (anonymous):

either way if its y you want, divide through by 15, if its x then divide by 9

OpenStudy (anonymous):

divide what by 15 or 9

OpenStudy (anonymous):

both sides, remember whatever you do to one side you have to do to the other

OpenStudy (anonymous):

so to get y, if we divide by 15 we get y = 9x/15

OpenStudy (anonymous):

0.6

OpenStudy (anonymous):

well we dont know what x is, so yeah it could be y = 0.6x but normally we like to leave it with whole numbers. y = 9x/15 it depends what your teacher wants you to do

OpenStudy (anonymous):

for x I got 0.6

OpenStudy (anonymous):

oh? whats y?

OpenStudy (anonymous):

y=0.6

OpenStudy (anonymous):

ahh did they give you that at the start of the question?

OpenStudy (anonymous):

x=1.6

OpenStudy (anonymous):

no gave us nothing for the question

OpenStudy (anonymous):

then how do you know y is 0.6 and not say... 10?

OpenStudy (anonymous):

or 100, or 123123, or 2341234.2134 or anything

OpenStudy (anonymous):

i'm confused

OpenStudy (anonymous):

well you have \[y = \frac{9x}{15}\]

OpenStudy (anonymous):

which is \[y = 0.6x\]

OpenStudy (anonymous):

if you decide to divide 9 by 15

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so it would be 1.6

OpenStudy (anonymous):

no. my point is you cant tell what x or y are.

OpenStudy (anonymous):

they can be any number. for instance what if x = 1? then we have y = 0.6(1) so y = 0.6

OpenStudy (anonymous):

but what if x = 2? y = 0.6(2) so y now = 1.2

OpenStudy (anonymous):

so there is no definite answer to find x or y

OpenStudy (anonymous):

no you need more information

OpenStudy (anonymous):

how about this questions how many terms are in this equation. (x+y)(xsquare+2xy+ysquare). Then work out the multiplication and combine like terms

OpenStudy (anonymous):

now there are times when you might have an equation like say y = 3x + 2 and the question would be "x is the distance, and y is the hight. Find out how high Hailey jumps when she travels 5 meters" then you could put "5" for x: y = 3(5) + 2 y = 17

OpenStudy (anonymous):

can you help me with the question above?